Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis Mackinaw Inc. processes a base chemical...

  1. Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

    Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 70,000 units of product were as follows:

    Standard Costs Actual Costs
    Direct materials 238,000 lbs. at $5.30 235,600 lbs. at $5.20
    Direct labor 17,500 hrs. at $16.40 17,900 hrs. at $16.80
    Factory overhead Rates per direct labor hr.,
    based on 100% of normal
    capacity of 18,260 direct
    labor hrs.:
    Variable cost, $4.10 $71,030 variable cost
    Fixed cost, $6.50 $118,690 fixed cost

    Each unit requires 0.25 hour of direct labor.

    Required:

    a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Material Price Variance $
    Direct Materials Quantity Variance $
    Total Direct Materials Cost Variance $

    b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Labor Rate Variance $
    Direct Labor Time Variance $
    Total Direct Labor Cost Variance $

    c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Variable factory overhead controllable variance $
    Fixed factory overhead volume variance $
    Total factory overhead cost variance $

Check My Work
Answer #1
a Material price variance=Actual quantity of material used*(Actual rate of material-Standard rate of material)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Actual material used=235600 lbs.
Standard rate of material=$ 5.30 per pound
Actual rate of material=$ 5.20 per pound
Material price variance=235600*(5.30-5.20)=23560=$ 23560 (Unfavorable)
Material Quantity variance=Standard rate of material*(Actual quantity consumed-Standard quantity)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Standard rate of material=$ 5.30 per pound
Standard quantity=238000 lbs.
Actual material used=235600 lbs.
Material quantity variance=5.30*(235600-238000)=-12720=$ 12720 (Favorable)
Total material cost variance=Material price variance+Material quantity variance
If the answer is negative, variance is favorable.Otherwise unfavorable.
Total material cost variance=23560+-12720=10840=$ 10840 (Unfavorable)
b Labor rate variance=Actual hours worked*(Actual rate of labor-Standard rate of labor)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Actual hours worked=17900 hours
Standard rate of labor=$ 16.40 per hour
Actual rate of labor=$ 16.80 per hour
Labor rate variance=17900*(16.80-16.40)=7160=$ 7160 (Unfavorable)
Labor time variance=Standard rate of labor*(Actual hours worked-Standard hours)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Standard rate of labor=$ 16.40 per hour
Standard hours=17500 hours
Actual hours worked=17900 hours
Direct labor time variance=16.40*(17900-17500)=6560=$ 6560 (Unfavorable)
Total labor cost variance=Labor rate variance+labor time variance
If the answer is negative, variance is favorable.Otherwise unfavorable.
Total labor cost variance=7160+6560=13720=$ 13720 (Unfavorable)
c Total factory overhead cost budgeted:
$
Variable cost (18260*4.10) 74866
Fixed cost (18260*6.50) 118690
Total 193556
Fixed factory overhead budgeted=18260*6.50=$ 118690
Variable factory overhead budgeted=$ 74866
Variable factory overhead controllable variance=Actual variable factory overhead-Budgeted variable factory overhead based on standard hours allowed
If the answer is negative, variance is favorable.Otherwise unfavorable.
Variable factory overhead controllable variance=71030-(17500*4.10)=71030-71750=-720=$ 720 (Favorable)
Fixed factory overhead volume variance=Absorbed fixed overhead-Budgeted fixed overhead
If the answer is negative, variance is unfavorable.Otherwise favorable.
Absorbed fixed overhead=Actual hours*rate per hour=17900*6.50=$ 116350
Fixed factory overhead volume variance=116350-118690=-2340=$ 2340 (Unfavorable)
Total factory overhead cost variance=Budgeted factory overhead-Actual factory overhead
If the answer is negative, variance is unfavorable.Otherwise favorable.
Total factory overhead cost variance=193556-(71030+118690)=193556-189720=3836=$ 3836 (Favorable)